OpenStudy (anonymous):

integrate (7x^2 + 5)^-1

OpenStudy (abb0t):

\(\large\int\limits \frac{1}{7x^2+5}dx\) is that correct?

OpenStudy (anonymous):

yup.

OpenStudy (abb0t):

play around with it to get it in the form of artan where you have \(ax^2+1\) which means factor out the 5.

OpenStudy (abb0t):

can you try that first :) if you need help, let me know.

OpenStudy (anonymous):

OMG thanks :P

OpenStudy (anonymous):

but does then does it become 1/a tan^-1(x/a) or something?? Sorry, i've forgotten the general formula.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (abb0t):

\[\int\limits \frac{1}{ax^2+1} = \arctan(x)\]

OpenStudy (anonymous):

I mean, what is the integral of \[\frac{ 1 }{ ax^2 + 1 }\]

OpenStudy (anonymous):

where does the a go??

OpenStudy (abb0t):

the "a" is the number in front of \(x^2\)!!

terenzreignz (terenzreignz):

Well, you could let \(u = x\sqrt{a}\) And work from there.

OpenStudy (anonymous):

Hold on guys, the derivative of arctan is \[\frac{ 1 }{ x^2 + 1 }\], so my question is that the a is a coefficient, but you need to multiply it or divide it to the arctan when you integrate it, so what do we do with it??

terenzreignz (terenzreignz):

Well... \[\large \int \frac{dx}{ax^2+1}\] \[let \ u = x\sqrt a\\du = \sqrt a \ dx\] \[\large = \frac1{\sqrt a}\int \frac{du}{u^2 +1}\] And work from there.

OpenStudy (anonymous):

ohh. ok, lemme do that.

OpenStudy (anonymous):

 This can be easily derived using the substitution\[x = a \tan \theta \] Here 'a' is a constant... :-)